OpenStudy (rainbow_dashie):

Let A = 2, B = , and C = . Find the value of each expression below. A × C + B A × B + C A + B - C A - B + C

OpenStudy (anonymous):

B = ? And C = ??

OpenStudy (anonymous):

What are the values of B and C given to you, you have not mentioned them in your question.. :)

OpenStudy (rainbow_dashie):

\[b=\frac{ 2 }{ 5 } c=\frac{ 3 }{ 8 }\]

OpenStudy (anonymous):

You tried it or not?

OpenStudy (anonymous):

I'll start the first one out: 2 x 3/8 x 2/5 = ?

OpenStudy (anonymous):

\[A \times C + B \\ 2 \times \frac{3}{8} + \frac{2}{5}\]

OpenStudy (anonymous):

You can start with first one, but you should start correct.. :P

OpenStudy (anonymous):

A X C Plus B ... It is PLUS SIGN in between.. :)

OpenStudy (anonymous):

Oh, yeah. My bad. Thanks @waterineyes

OpenStudy (rainbow_dashie):

\[2\times \frac{ 2 }{ 5 }+\frac{ 3 }{ 8 }\]

OpenStudy (anonymous):

Yup.

OpenStudy (rainbow_dashie):

my computer is crap its slow sorry

OpenStudy (anonymous):

@Rainbow_dashie , is that first one or second one?

OpenStudy (anonymous):

\[2\times \frac{ 2 }{ 5 }+\frac{ 3 }{ 8 }\]

OpenStudy (rainbow_dashie):

2*0.4+0.375

OpenStudy (rainbow_dashie):

\[2\times0.4+0.375\]

OpenStudy (anonymous):

you are taking them into decimals?? you should go with LCM or making common denominator approach.. Do you know something about these?

OpenStudy (rainbow_dashie):

answer my questions said to do tht

OpenStudy (anonymous):

Hehe...

OpenStudy (anonymous):

Hi?

OpenStudy (anonymous):

Write his username properly.. :P

OpenStudy (rainbow_dashie):

i'm sorry AnswerMyQuestions there that better

OpenStudy (anonymous):

Answermyquestions , like this.. Otherwise, I at first time had difficulty in understanding : Now you read : water in eyes said answer my questions, and Rainbow got dashie.. :P

OpenStudy (rainbow_dashie):

ok were getting off topic math is the thing im after

OpenStudy (anonymous):

See, if you think you can handle decimals, then yes you can do that.. Now multiply 2 with 0.4 and add 0.375 in it.. :)

OpenStudy (rainbow_dashie):

\[1.175\]

OpenStudy (rainbow_dashie):

@waterineyes

OpenStudy (anonymous):

Good.. :)

OpenStudy (anonymous):

Are you given with some options??

OpenStudy (rainbow_dashie):

My computer died I had to revive it

OpenStudy (anonymous):

If you have water with you, then yes you can.. :P

OpenStudy (rainbow_dashie):

Ok this question is no longer in use my study island reset